OpenStudy (cathyangs):

If z=a+bi for real numbers a and b and z^2=21-201, what is the value of |a|+|b|?

OpenStudy (earthcitizen):

z=a+jb

OpenStudy (earthcitizen):

you can use de moivre's theorem

OpenStudy (cathyangs):

agg, sorry it's actually z^2=21-20i

OpenStudy (earthcitizen):

\[z=(-b \pm \sqrt{b^2 - 4ac})/2a\]

OpenStudy (earthcitizen):

yh, re-arrange to give z^2+i20-21

OpenStudy (jamesj):

actually note that (a+bi)^2 = (a^2 - b^2) + 2abi. Now set that equal to 21-20i and solve for a and b.

OpenStudy (earthcitizen):

use the almighty formula , identify a, b and c and subs.t into the formula

OpenStudy (earthcitizen):

@ JamesJ you do realize that that is a confusing method as per not related to complex numbers

OpenStudy (cathyangs):

ah...thanks!

OpenStudy (jamesj):

What do you mean? It's absolutely germane to complex numbers.

OpenStudy (jamesj):

In fact, try and use your method and you'll get stuck.

OpenStudy (earthcitizen):

interesting, i might have not come across it

OpenStudy (anonymous):

\[z=a+bi \implies z^2=(a+bi)^2=a^2+2abi-b^2=21-20i \implies (a^2-b^2)=21\] \[2ab=-20\] Now solve. Two equations two unknowns.

OpenStudy (anonymous):

That first line should read: \[a^2-b^2=21^*\]

OpenStudy (cathyangs):

wait, JamesJ, the a and b formula you put up there doesn't work...or am i wrong?

OpenStudy (anonymous):

As James said, don't use QF. lolz

OpenStudy (cathyangs):

oh...never mind! :P

OpenStudy (jamesj):

mal19 has laid it out for you. Two complex numbers are equal if and only if their real and imaginary components are equal. Hence if z = a+bi and z^2 = 21 - 20i, it must be that: \[ a^2 - b^2 = 21 \ \ \ \ 2ab = -21 \]

OpenStudy (anonymous):

Yup :)

OpenStudy (jamesj):

**second equation should be = -20

OpenStudy (cathyangs):

but how do I get to the absolute value?

OpenStudy (jamesj):

Solve for a and b. Then take the absolute value at the very end.

OpenStudy (cathyangs):

... I got something that's definitely not an answer(which is going to be between 3 and 7, inclusive)

OpenStudy (jamesj):

Using the second equation, 2ab = -20, then b = -10/a. Substitute that into the equation a^2 - b^2 = 21 and solve for a. Take those solutions and find the corresponding b.

OpenStudy (anonymous):

So we have: \[2ab=-20 \implies b=\frac{-10}{a}\] But: \[a^2-b^2=21\] as well so: \[a^2- \left( \frac{-10}{a} \right)^2=21 \implies a^2-\frac{100}{a^2}=21\] Multiply by a^2 to give: \[a^4-100=21a^2 \implies a^4-21a^2-100=0\] Make the substitution y=a^2 and we get: \[y^2-21y-100=0 \implies (y-25)(y+4)=0 \implies (a^2-25)(a^2+4)=0 \implies a= \pm 5\] \[ a= \pm 2i\] Unless I'm doing something completely wrong, lol. Now solve b?

OpenStudy (anonymous):

after the implies it should be: \[a= \pm 5\]

OpenStudy (jamesj):

Right. a must be real, so the values of a are +5 or -5. Find now the corresponding values of b.

OpenStudy (anonymous):

And I'm assuming you only care about the real answer (or else the number wouldn't be in the form a+bi (think, if a=2i then you have (2+b)i, which is different) So a=5 (its an abs value so you don't need to worry about sign)

OpenStudy (anonymous):

Exactly^

OpenStudy (cathyangs):

the answer is 7, right? thanks everyone!!! (If I could medal multiple times, I would) can you guys stick around? I have other questions. I swear I'm seriously stuck on these (I'm not lazy) I'd really like to know how to solve them,(our teacher gave them as "extra practice" over the break, but he's not talking about then in class)

OpenStudy (anonymous):

Correct, if my mental arithmetic is correct.

OpenStudy (jamesj):

By the way, the two solutions 5 - 2i and -5 + 2i are additive opposites, which is exactly what we would expect. You can and should verify that they do indeed square to 21 - 20i.

OpenStudy (jamesj):

Post the new question on the left, thanks.

OpenStudy (earthcitizen):

i guess we learn something new everyday. I'm quite sure my method would give the same answer

OpenStudy (earthcitizen):

i can show you and i won't get stuck ^_^

OpenStudy (earthcitizen):

\[z^2=21-i20=>z=\sqrt{21-i20}\]

OpenStudy (earthcitizen):

\[z=(21-i20)^{1/2}\]

OpenStudy (anonymous):

You aren't solving z...You're solving a and b... Now explain how to find the sqrt of that number as well please :)

OpenStudy (earthcitizen):

yh, can't we find the two solutions of z and then find the magnitude of a and b from those two solutions ? for ex. if z1=a1+ib1 and z2=a2+ib2

OpenStudy (jamesj):

right, exactly. Where do you go from there.

OpenStudy (earthcitizen):

we will need to transform that into polar form

OpenStudy (earthcitizen):

\[r<\theta\]

OpenStudy (anonymous):

Given an arbitrary complex number z and w multiplying them would give: \[z*w=rq(\cos(\theta+\phi)+isin(\theta+\phi))\] Assuming a modulus of r and q for z and w respectively. How would that help?

OpenStudy (anonymous):

Also, my steps (way) above make the assumption that: \[a \ne 0\] (I believe we have to make that assumption***)

OpenStudy (earthcitizen):

yh, i meant..

OpenStudy (earthcitizen):

\[z ^{-n}=r ^{-n}(\cos(n \theta)-isin(n \theta))\]

OpenStudy (earthcitizen):

\[z=(21-i20)^{1/2}=(29<44)^{1/2}\]

OpenStudy (earthcitizen):

\[z=(29^{1/2})<44/2)\]

OpenStudy (anonymous):

What are you doing? lol

OpenStudy (earthcitizen):

i'm trying to find the two roots to the equation

OpenStudy (earthcitizen):

lol

OpenStudy (anonymous):

Okay, go for it :P

OpenStudy (earthcitizen):

you said you me to solve z=(21-j20)^1/2

OpenStudy (earthcitizen):

you said you me to solve z=(21-j20)^1/2\[z=(21-j20)^{1/2}\]

OpenStudy (earthcitizen):

i do know \[z _{1}=5.4<-22^{o}, z _{2}=5.4<158^{o}\]

OpenStudy (jamesj):

Degrees! You should carry the arctan(-20/21) and then work out the cos and sin of half that quantity (+ pi) so you recover the exact solutions.

OpenStudy (earthcitizen):

\[\therefore z _{1}=5.4(\cos(-22)+jsin(-22)) = 4.6-j2\]

OpenStudy (earthcitizen):

yh, you can't add pi if x>0

OpenStudy (earthcitizen):

\[z _{2} = -5+j2\]

OpenStudy (earthcitizen):

as you can see we both arrived at the same ans